OpenStudy (anonymous):

Use properties of logarithms to condense the logarithmic expression. Write the expression as a single logarithm whose coefficient is 1. Where possible, evaluate logarithmic expressions without using a calculator. log_3(567)-log_3(7)

OpenStudy (phi):

you could use log(a) - log(b) = log(a/b)

OpenStudy (phi):

as a start

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- notice the 2nd rule listed there

OpenStudy (anonymous):

log_3(567^3/7^3) ?

OpenStudy (phi):

no ^3 just 567/7

OpenStudy (anonymous):

Apparently that answer is missing inputs

OpenStudy (anonymous):

I'm not sure what that even means

OpenStudy (phi):

we are not done yet

OpenStudy (anonymous):

(this homework is online)

OpenStudy (phi):

log_3(567)-log_3(7) = ?

OpenStudy (phi):

use the rule log(a) - log(b)= log(a/b)

OpenStudy (anonymous):

right so what would you do from log_3(81)? 81^X=3?

OpenStudy (anonymous):

or wait 3^x=81? then solve?

OpenStudy (phi):

yes, you get log_3(81) if you can write 81 as 3^x (and we can) we can simplify some more

OpenStudy (anonymous):

so is the answer what x is, which is 4?

OpenStudy (phi):

we replace 81 with 3^4 in log_3(81) log_3(3^4) now use the rule log(a^b) = b log(a)

OpenStudy (phi):

in other words log_3(81)= 4 log_3(3) last rule (finally!) log_3(3) = 1

OpenStudy (anonymous):

4 ended up being the correct answer

OpenStudy (phi):

yes 4 log_3(3) = 4*1 = 4 so we found log_3(567)-log_3(7) = 4

OpenStudy (anonymous):

oh okay thats where the one comes from I get it now!

OpenStudy (anonymous):

That one makes sense to me. How about this one? Use properties of logarithms to condense the logarithmic expression. Write the expression as a single logarithm whose coefficient is 1. Where possible, evaluate logarithmic expressions without using a calculator. 8ln(x+14)-5ln(x)

OpenStudy (phi):

you can "bring in" the 8 (and the 5) using a log(b)= log(b^a)

OpenStudy (anonymous):

so log(x+14)^8?

OpenStudy (phi):

yes, but it's more clear to write it log( (x+14)^8 ) because the 8 is inside the log

OpenStudy (phi):

do the same for 5 log(x)

OpenStudy (anonymous):

okay so I have log((x+14)^8)-log((x^5)). now what?

OpenStudy (phi):

now use log(a) - log(b) = log(a/b) and we should use ln (which is short of log base e) to match the problem

OpenStudy (anonymous):

Can you show me how to do that with this one I'm confused lol

OpenStudy (phi):

log( (hairy expression)) - log((messy stuff)) = log( (hairy expression)/(messy stuff) )

OpenStudy (anonymous):

log((x+14)^8/(x^5)) ?

OpenStudy (phi):

yes. Just to be clear, it is \[\ln \left( \frac{ (x+14)^8 }{ x^5 } \right)\]

OpenStudy (anonymous):

Thank you. I missed class so I am unaware of all the properties. Can you please help with one more:)

OpenStudy (anonymous):

Use properties of logarithms to condense the logarithmic expression. Write the expression as a single logarithm whose coefficient is 1. Where possible, evaluate logarithmic expressions without using a calculator.

OpenStudy (anonymous):

I attatched it as a file

OpenStudy (anonymous):

I am suppose to fill in the blank after log_7

OpenStudy (phi):

do it step by step first, "condense" log_7(x) - log_7(y)

OpenStudy (anonymous):

log_7(x/y)

OpenStudy (phi):

now we have (1/7) log_7(x/y) use a log(b) = log(b^a) in this case, b is (x/y)

OpenStudy (anonymous):

log((x/7)^1/7)

OpenStudy (anonymous):

x/y***

OpenStudy (phi):

yes , and base 7 i.e. log_7

OpenStudy (phi):

\[\log_7\left( \left( \frac{ x }{ y } \right)^{\frac{ 1 }{ 7 }} \right)+5 \log_7\left( x-6\right)\]

OpenStudy (phi):

now condense the 2nd term

OpenStudy (phi):

using a log(b) = log(b^a)

OpenStudy (anonymous):

log_7((x-6)^5)

OpenStudy (phi):

ok, last step log_7((x/y)^1/7) + log_7( (x-6)^5) use log(a)+log(b) = log(a b) (i.e multiply a times b)

OpenStudy (anonymous):

log_7((x/y)^1/7 * (x-6)^5)

OpenStudy (anonymous):

or is there a way to actually simplify that

OpenStudy (phi):

no. that is what it is. A trifle ugly but we mustn't be rude to its face.

OpenStudy (anonymous):

haha okay so that's the answer?

OpenStudy (phi):

yes, that is the answer. if you want more background on this, you can watch http://www.khanacademy.org/math/algebra2/logarithms-tutorial/logarithm_properties/v/introduction-to-logarithm-properties

OpenStudy (anonymous):

thank you very much. You have been a huge help. I have more questions but I must be waring you out, should I repost?